Given a population consists of three numbers (2,5,9). Consider all possible samples of size 2 which can be drawn from the population. Find the standard deviation of the sampling distribution of the sample means.


1
Expert's answer
2021-11-16T14:15:09-0500

a.

 We have population values "2,5,9" population size "N=3"

"\mu=\dfrac{2+5+9}{3}=\dfrac{16}{3}""\sigma^2=\dfrac{1}{3}((2-16/3)^2+(5-16/3)^2+(9-16/3)^2"

"=\dfrac{74}{9}"




"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{74}{9}}=\dfrac{\sqrt{74}}{3}"


We have population values "2,5,9" population size "N=3" and sample size "n=2." Thus, the number of possible samples which can be drawn without replacement is



"\dbinom{N}{n}=\dbinom{3}{2}=3""\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 2,5 & 7/2 \\
 \hdashline
 2 & 2,9 & 11/2 \\
 \hdashline
 3 & 5,9 & 7 \\
 \hline
\end{array}"

The sampling distribution of the sample means.



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
& \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 & 7/2 & 1 & 1/3 & 7/6 & 49/12 \\
 \hdashline
 & 11/2 & 1 & 1/3 & 11/6 & 121/12 \\
 \hdashline
 & 7 & 1 & 1/3 & 7/3 & 49/3 \\
 \hdashline
 Total & & 3 & 1 & 16/3 & 183/6 \\ \hline
\end{array}"




"E(\bar{X})=\sum\bar{X}f(\bar{X})=\dfrac{16}{3}"


The mean of the sampling distribution of the sample means is equal to the the mean of the population.


"E(\bar{X})=\mu_{\bar{X}}=\dfrac{16}{3}=\mu"




"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"




"=\dfrac{183}{6}-(\dfrac{16}{3})^2=\dfrac{37}{18}"




"\sigma_{\bar{X}}=\sqrt{Var(\bar{X})}=\sqrt{\dfrac{37}{18}}=\dfrac{\sqrt{74}}{6}"

Verification:


"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{\dfrac{74}{9}}{2}(\dfrac{3-2}{3-1})"




"=\dfrac{37}{18}, True""\bar{X}\sim N(\dfrac{16}{3}, \dfrac{37}{18})"